OpenStudy (anonymous):

HELP PLEASE!!!! Which of the function(s) have a horizontal asymptote of y = 1? (Select all that apply.) I. y=(x^2+1)/(x^2-1) II. y=(x-1)/(x^2+1) III. y=(x^2-1)/(x^2+1)

OpenStudy (anonymous):

@xmarioz25 help please?

OpenStudy (anonymous):

Divide by the highest exponent . Apply the limit as x tends to infinity 1/x^r = 0

OpenStudy (anonymous):

It'd be the first and third one.

OpenStudy (anonymous):

yeah i had the third one down but i didnt know about the first one.

OpenStudy (anonymous):

\[1. y = \frac{ \frac{x^2 }{x^2} + \frac{1}{x^2} }{ (x^2/x^2 -1/x^2) }\] \[y = (1 + 0 )/(1-0) = 1\]

OpenStudy (anonymous):

makes perfect sense thanks!